# Pretraining with Random Noise for Fast and Robust Learning without Weight Transport

### v1

@@ Abstract
The brain prepares for learning even before interacting with the environment, by refining and optimizing its structures through spontaneous neural activity that resembles random noise. However, the mechanism of such a process has yet to be thoroughly understood, and it is unclear whether this process can benefit the algorithm of machine learning. Here, we study this issue using a neural network with a feedback alignment algorithm, demonstrating that pretraining neural networks with random noise increases the learning efficiency as well as generalization abilities without weight transport. First, we found that random noise training modifies forward weights to match backward synaptic feedback, which is necessary for teaching errors by feedback alignment. As a result, a network with pre-aligned weights learns notably faster than a network without random noise training, even reaching a convergence speed comparable to that of a backpropagation algorithm. Sequential training with both random noise and data brings weights closer to synaptic feedback than training solely with data, enabling more precise credit assignment and faster learning. We also found that each readout probability approaches the chance level and that the effective dimensionality of weights decreases in a network pretrained with random noise. This pre-regularization allows the network to learn simple solutions of a low rank, reducing the generalization loss during subsequent training. This also enables the network robustly to generalize a novel, out-of-distribution dataset. Lastly, we confirmed that random noise pretraining reduces the amount of meta-loss, enhancing the network ability to adapt to various tasks. Overall, our results suggest that random noise training with feedback alignment offers a straightforward yet effective method of pretraining that facilitates quick and reliable learning without weight transport.
@@ 1 Introduction
The brain refines its network structure and synaptic connections even before birth, without exposure to sensory stimuli [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. In the early developmental stages, the spontaneous neuronal activity that appears in various brain regions is considered to play a critical role during the development of neuronal circuits by pruning neural wirings and adjusting synaptic plasticity [27, 28, 29, 30, 31]. If this activity is disrupted during the developmental stages, the outcome can be long-lasting neuronal deficits. Computational studies suggest that such a refined network structure enables certain crucial functions of the brain, such as initializing function and efficient learning [32, 33, 34, 35]. These experimental and theoretical studies commonly indicate that spontaneous, random neuronal activity plays a critical role in the development of the biological neural network before data are encountered by the network. However, the detailed mechanism of how these prenatal processes contribute to learning after birth, i.e., with subsequent sensory stimuli, remains elusive.
@@
At the synaptic level, learning can be defined as the process by which the brain adjusts the strength of synaptic connections between neurons to optimize the network for a specific task. The synaptic weights of each neuron can change to minimize the error between the expected and the actual output of a task, often referred to as the credit assignment problem [36]. However, in general, it is not well known how individual neurons modify these synaptic connections and thus achieve a network goal under a condition in which numerous neurons are linked in multiple layers. In other words, how neurons can estimate errors to modify their synaptic connections during learning remains unknown.
@@
In machine learning, backpropagation algorithms have successfully addressed this issue – even in deep neural networks [37, 38, 39]. Backpropagation can provide feedback with regard to forward errors through the symmetric copying of forward weights via a backward process. During this process, a structural constraint, i.e., symmetric forward and backward weights, is necessary to assign proper error values to individual neurons [40, 41]. However, this process appears to be biologically implausible due to the weight transport problem [42, 43, 44, 36], in which individual neurons must somehow be aware of the exact synaptic connections of their downstream layers to update their weights, a state considered to be practically impossible in a biological brain.
@@
An alternative algorithm, feedback alignment, achieves successful network training even without weight transport by employing fixed random feedback pathways [44]. Studies in relation to this show that a network can align its weights to synaptic feedback during data training, and this simple process enables error backpropagation. It has been shown that soft alignment between forward weights and synaptic feedback, which can be achieved during learning with massive amounts of data, is enough to back-propagate errors. This finding may provide a biologically plausible scenario in which the credit assignment problem can be resolved, yet there is an issue remaining — the process requires massive data learning to develop the structural constraint. This cannot be addressed even with currently known advanced learning rules [45, 46].
@@
This situation is contradictory to the notion that the brain can learn even with very limited experience in the initial stages of life [47, 48, 49, 50]. Thus, the question arises as to how early brains can estimate and assign errors for learning with limited experience. To address this issue, here we focus on the role of spontaneous activity at the prenatal stage in the brain, showing that training random noise, which mimics spontaneous random activity in prenatal brains, is a possible solution; random noise training aligns the forward weights to synaptic feedback, enabling precise credit assignment and fast learning. We also observed that random noise training can pre-regulate the weights and enable robust generalization. Our findings suggest that random noise training is a core mechanism of prenatal learning in biological brains and that it may provide a simple algorithm for the preconditioning of artificial neural networks for fast and robust learning without the weight transport process.
@@ 2 Preliminaries
Biological and artificial neural networks have different structures and functionalities, but they share certain factors in common, such that information is processed through multiple hierarchical layers of neurons with a nonlinear response function. In the current study, we consider a multi-layer feedforward neural network for pattern classification, $f_{\theta}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{d}$, parameterized by $\theta=\{\mathbf{W}_{l},\mathbf{b}_{l}\}_{l=0}^{L-1}$. It takes input $\mathbf{x}\in\mathbb{R}^{m}$ and outputs a vector $\mathbf{y}\in\mathbb{R}^{d}$ with $L$ layers. Through a forward pass, the network computes a hidden layer output by propagating the input through the network layers, as follows:
@@
, where $\mathbf{W}_{l}$ is the forward matrix, $\mathbf{b}_{l}$ is the bias vector, and $\phi$ is the nonlinear activation function. In the first layer $l=0$, $\mathbf{h}_{l}=\mathbf{x}$. We used a rectified linear unit (ReLU) activation function, $\phi(x)=\max(0,x)$. In the last layer $l={L-1}$, we used a softmax function, $\phi_{y}(x)=\text{softmax}(x)=\{e^{x_{i}}/\sum_{j=1}^{d}e^{x_{j}}\}_{i=1}^{d}$. Thus, the network outputs a probability distribution over $d$ classes. After the forward pass, the amount of error is calculated by measuring the difference between the network output $f_{\theta}(\mathbf{x})$ and the target label $\mathbf{y}$. We used the cross-entropy loss [51], which is defined as follows:
$\mathcal{L}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{d}y_{ij}\log f_{\theta}(\mathbf{x}_{i})_{j}$ (2)
, where $N$ is the number of samples, $d$ is the number of classes, and $y_{ij}$ is the target label for the $i$-th sample and the $j$-th class. The purpose of learning is to minimize the error $\mathcal{L}(\theta)$. To achieve this, the network parameters $\theta$ are adjusted by assigning credit to the weights that contribute to the error, which is known as the credit assignment problem.
@@ 2.1 Backpropagation and weight transport problem
To solve the credit assignment problem, backpropagation [37] computes the gradient of errors with respect to the weights and uses it as a teaching signal to modulate the aforementioned parameters. The gradient is calculated by the chain rule, with propagation from the output layer to the input layer, as follows:
@@
,where $\eta$ is the learning rate. The backpropagation algorithm successfully solves the credit assignment problem, but it requires heavy computation to use the complete information of the synaptic weights of the next layer to update the current weights. Notably, backpropagation is considered as biologically implausible, because it is impossible, in the brain, to transmit the synaptic weights from the next layer to the current layer. This is known as the weight transport problem.
@@ 2.2 Feedback alignment
To address the weight transport problem, the idea of feedback alignment [44] was proposed as a biologically plausible alternative to backpropagation. In feedback alignment, the backward synaptic feedback is replaced with a random, fixed weight matrix $\mathbf{B}_{l}$ in the feedback path, as follows:
@@
The only difference between backpropagation and feedback alignment is the replacement of the transpose of the forward matrix $\mathbf{W}_{l}$ with the fixed random feedback matrix $\mathbf{B}_{l}$ to calculate the error signal. The fact that the network can learn tasks from error teaching signals that are calculated from random feedback is explained by the observation that the network modifies the forward matrices $\mathbf{W}_{l}$ to match the transpose of the feedback matrices $\mathbf{B}_{l}$ roughly during training. This makes the error teaching signal (5) similar to backpropagation (3), thus enabling the network to learn the task.
3 Random noise training with feedback alignment
Algorithm 1 Random noise training
  Input: Network $f_{\theta}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{d}$, learning rate $\eta$, number of examples $N_{\text{examples}}$, batch size $N_{\text{batch}}$
  Output: Trained network $f_{\theta^{{}^{\prime}}}$
  for $i=1$ to $N_{\text{examples}}//N_{\text{batch}}$ do
     x from Gaussian distribution $\mathcal{N}(0,I)$, y from discrete uniform distribution $\mathcal{U}(0,d-1)$
     Forward pass: $\hat{y}=f_{\theta}(x)$
     Calculate loss: $\mathcal{L}=\text{cross\_entropy}(\hat{y},y)$
     Backward pass: $\forall l$, $\mathbf{W}_{l}=\mathbf{W}_{l}-\eta\delta_{l}\mathbf{h}_{l}^{T}$
  end for
@@ 3 Random noise training with feedback alignment
During the developmental stage, spontaneous neural activity in the brain plays a critical role in shaping and refining neural circuits. Initially wired immature neural circuits undergo modifications of their connections through the processes of regulated cell formation, apoptosis, and synapse refinement through spontaneous neural activity [27, 28, 29, 30, 31]. These pre-sensory activities and development processes are universal across sensory modalities, such as the visual, auditory, and motor systems. We focus here on a few characteristics of spontaneous neural activity in the brain. Spontaneous neural activity is not correlated to external stimuli but can refine and optimize neural circuits, before interaction with the external world can take place.
@@
Here, we propose a type of random training that is inspired by the spontaneous and prenatal neural activity in the brain to pretrain the neural network (Algorithm 1). In every batch, we sampled random noise inputs $\mathbf{x}$ from a Gaussian distribution $\mathcal{N}(0,I)$ and random labels $\mathbf{y}$ from a discrete uniform distribution $\mathcal{U}(0,N_{\text{readout}}-1)$, without any correlation. The network $f_{\theta}$ was initialized with random weights and trained with the feedback alignment algorithm. In this study, we examined the effects of random noise training on the subsequent learning processes in model neural networks to understand the potential benefits of pretraining with random noise in biological brains and whether this strategy is applicable to machine learning algorithms.
@@ 4.1 Weight alignment to synaptic feedback during random noise training
To simulate a neural network initially wired by random weights and fixed random synaptic feedback, we adopted a network setting from the feedback alignment algorithm (Figure 1a) in which the weight transport problem can be avoided through the use of fixed random synaptic feedback. Thus, unlike backpropagation, this process is considered possibly to exist in biological neural networks with local synaptic connections (Figure 1b). We used a two-layer feedforward neural network with ReLU nonlinearity for classification, $f_{\theta}:\mathbb{R}^{784}\rightarrow\mathbb{R}^{10}$ with 100 neurons in the hidden layer. By means of random noise training (Algorithm 1), we trained the neural network with random inputs sampled from a Gaussian distribution $\mathcal{N}(0,I)$, with labels also randomly sampled independently (Figure 1c).
@@
We observed that the loss decreased noticeably during random training, even in the absence of meaningful data and even when $\mathbf{x}$ and $\mathbf{y}$ simply are randomly paired (Figure 1d). During the random training process, we focused on the alignment between the forward weights and the synaptic feedback. As described in the literature [44], the alignment of $\mathbf{W}_{l}$ and $\mathbf{B}_{l}$, i.e., similarity between $\delta_{\text{BP}}$ and $\delta_{\text{FA}}$, is crucial for calculating the error teaching signal precisely. To evaluate the alignment, we used cosine similarity, which is widely used for measuring the distance between two vectors.
@@
Definition. Given the forward weights $\mathbf{W}_{l}\in\mathbb{R}^{m\times n}$ and backward weights $\mathbf{B}_{l}\in\mathbb{R}^{n\times n}$, we used the cosine angle as a measure of the alignment of individual neurons. We claim that $\mathbf{W}_{l}$ and $\mathbf{B}_{l}$ are aligned if $\mathbb{E}(\theta_{i})=\mathbb{E}(\angle(\mathbf{W}_{l}^{T})_{i},(\mathbf{B}_{l})_{i})$ decreases asymptotically during training [44].
@@
Notably, we found that the weights of neurons are aligned to the corresponding synaptic feedback weights during the random training process (Figure 1e). We also observed that the angle between the forward weights and synaptic feedback of individual neurons in the hidden layer decreased asymptotically during random training. In a randomly initialized network, the alignment angle appeared to be close to 90°, demonstrating that the backward error signal is randomly distributed (Figure 1e, left, alignment angle in an untrained network vs. 90°, $n=100$, one-sample t-test, NS, $P=0.492$). However, after random training, the alignment angle decreased significantly, implying that the backward teaching signal becomes valid to back-propagate errors (Figure 1d, right, alignment angle in an untrained network vs. a randomly trained network, $n=100$, two-sample t-test, ${}^{*}P=<0.001$). We confirmed that this is not simply due to input bias under a particular condition but is reproduced robustly with various input conditions (Figure 1f). These results suggest that neural networks can pre-learn how to back-propagate errors through random noise training.

### v2

@@ Abstract
The brain prepares for learning even before interacting with the environment, by refining and optimizing its structures through spontaneous neural activity that resembles random noise. However, the mechanism of such a process has yet to be understood, and it is unclear whether this process can benefit the algorithm of machine learning. Here, we study this issue using a neural network with a feedback alignment algorithm, demonstrating that pretraining neural networks with random noise increases the learning efficiency as well as generalization abilities without weight transport. First, we found that random noise training modifies forward weights to match backward synaptic feedback, which is necessary for teaching errors by feedback alignment. As a result, a network with pre-aligned weights learns notably faster and reaches higher accuracy than a network without random noise training, even comparable to the backpropagation algorithm. We also found that the effective dimensionality of weights decreases in a network pretrained with random noise. This pre-regularization allows the network to learn simple solutions of a low rank, reducing the generalization error during subsequent training. This also enables the network to robustly generalize a novel, out-of-distribution dataset. Lastly, we confirmed that random noise pretraining reduces the amount of meta-loss, enhancing the network ability to adapt to various tasks. Overall, our results suggest that random noise training with feedback alignment offers a straightforward yet effective method of pretraining that facilitates quick and reliable learning without weight transport.
@@ 1 Introduction
The brain refines its network structure and synaptic connections even before birth, without exposure to sensory stimuli [1, 2, 3, 4]. In the early developmental stages, the spontaneous neuronal activity that appears in various brain regions is considered to play a critical role during the development of neuronal circuits by pruning neural wiring and adjusting synaptic plasticity [5, 6, 7, 8]. If this activity is disrupted during the developmental stages, the outcome can be long-lasting neuronal deficits [9, 10, 11]. Computational studies suggest that such a refined network structure enables certain crucial functions of the brain, such as initializing function [12, 13, 14, 15] and structure [16, 17]. These experimental and theoretical studies commonly indicate that spontaneous, random neuronal activity plays a critical role in the development of the biological neural network before data are encountered by the network. However, the detailed mechanism of how these prenatal processes contribute to learning after birth, i.e., with subsequent sensory stimuli, remains elusive.
@@
At the synaptic level, learning can be defined as the process by which the brain adjusts the strength of synaptic connections between neurons to optimize the network for a specific task [18, 19, 20]. The synaptic weights of each neuron can change to minimize the error between the expected and the actual output of a task, often referred to as the credit assignment problem [21, 22, 23]. However, in general, it is not well known how individual neurons modify these synaptic connections and thus achieve a network goal under a condition in which numerous neurons are linked in multiple layers. In other words, how neurons can estimate errors to modify their synaptic connections during learning remains unknown.
@@
In machine learning, backpropagation algorithms have successfully addressed this issue – even in deep neural networks [24, 25, 26]. Backpropagation can provide feedback with regard to forward errors through the symmetric copying of forward weights via a backward process. During this process, a structural constraint, i.e., symmetric forward and backward weights, is necessary to assign proper error values to individual neurons [27, 28, 29]. However, this process appears to be biologically implausible due to the weight transport problem [30, 31, 32, 23], in which individual neurons must somehow be aware of the exact synaptic connections of their downstream layers to update their weights, a state considered to be practically impossible in a biological brain.
@@
An alternative algorithm, feedback alignment, achieves successful network training even without weight transport by employing fixed random feedback pathways [32]. This study shows that a network can align its weights to synaptic feedback during data training, and this simple process enables error backpropagation. It has been shown that soft alignment between forward and backward weights, which can be achieved during learning data, is enough to back-propagate errors. This finding may provide a biologically plausible scenario in which the credit assignment problem can be resolved, yet there is an issue remaining — the process requires massive data learning to develop the structural constraint, and it significantly underperforms compared to backpropagation on challenging tasks [33, 34]. This cannot be addressed even with currently known advanced learning rules [35, 36, 37, 38, 39, 40].
@@
This situation is contradictory to the notion that the brain can learn even with very limited experience in the initial stages of life [41, 42, 43, 44]. Thus, the question arises as to how early brains can estimate and assign errors for learning with limited experience. To address this issue, here we focus on the role of spontaneous activity at the prenatal stage in the brain, showing that training random noise, which mimics spontaneous random activity in prenatal brains, is a possible solution; random noise training aligns the forward weights to synaptic feedback, enabling precise credit assignment and fast learning. We also observed that random noise training can pre-regulate the weights and enable robust generalization. Our findings suggest that random noise training may be a core mechanism of prenatal learning in biological brains and that it may provide a simple algorithm for the preconditioning of artificial neural networks for fast and robust learning without the weight transport process.
@@ 2 Preliminaries
Biological and artificial neural networks have different structures and functionalities, but they share certain factors in common, such that information is processed through hierarchical layers of neurons with a nonlinear response function. In the current study, we consider a multi-layer feedforward neural network for pattern classification, $f_{\theta}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{d}$, parameterized by $\theta=\{\mathbf{W}_{l},\mathbf{b}_{l}\}_{l=0}^{L-1}$. It takes input $\mathbf{x}\in\mathbb{R}^{m}$ and outputs a vector $\mathbf{y}\in\mathbb{R}^{d}$ with $L$ layers. Through a forward pass, the network computes a hidden layer output by propagating the input through the network layers, as follows:
@@
, where $\mathbf{W}_{l}$ is the forward weights, $\mathbf{b}_{l}$ is the bias vector, and $\phi$ is the nonlinear activation function. In the first layer $l=0$, $\mathbf{h}_{l}=\mathbf{x}$. We used a rectified linear unit (ReLU) activation function, $\phi(x)=\max(0,x)$. In the last layer $l={L-1}$, we used a softmax function, $\phi_{y}(x)=\text{softmax}(x)=\{e^{x_{i}}/\sum_{j=1}^{d}e^{x_{j}}\}_{i=1}^{d}$. Thus, the network outputs a probability distribution over $d$ classes. After the forward pass, the amount of error is calculated by measuring the difference between the network output $f_{\theta}(\mathbf{x})$ and the target label $\mathbf{y}$. We used the cross-entropy loss [45], which is defined as follows:
$\mathcal{L}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{d}\mathbf{y}_{ij}\log f_{\theta}(\mathbf{x}_{i})_{j}$ (2)
, where $N$ is the number of samples, $d$ is the number of classes, and $\mathbf{y}_{ij}$ is the target label for the $i$-th sample and the $j$-th class. The purpose of learning is to minimize the error $\mathcal{L}(\theta)$. To achieve this, the network parameters $\theta$ are adjusted by assigning credit to the weights that contribute to the error, which is known as the credit assignment problem.
@@ 2.1 Backpropagation and weight transport problem
To solve the credit assignment problem, backpropagation (BP) [24] computes the gradient of errors with respect to the weights and uses it as a teaching signal to modulate the aforementioned parameters. The gradient is calculated by the chain rule, with propagation from the output layer to the input layer, as follows:
@@
, where $\eta$ is the learning rate. The backpropagation algorithm successfully solves the credit assignment problem, but it requires heavy computation to use the complete information of the synaptic weights of the next layer to update the current weights. Notably, backpropagation is considered as biologically implausible, because it is impossible, in the brain, to transmit the synaptic weights from the next layer to the current layer. This is known as the weight transport problem [30, 31].
@@ 2.2 Feedback alignment
To address the weight transport problem, the idea of feedback alignment (FA) [32] was proposed as a biologically plausible alternative to backpropagation. In feedback alignment, the backward synaptic feedback is replaced with a random, fixed weight matrix $\mathbf{B}_{l}$ in the feedback path, as follows:
@@
The only difference between backpropagation and feedback alignment is the replacement of the transpose of the forward weights $\mathbf{W}_{l}$ with the fixed random feedback weights $\mathbf{B}_{l}$ to calculate the error signal. The fact that the network can learn tasks from error teaching signals that are calculated from random feedback is explained by the observation that the network modifies the forward weights $\mathbf{W}_{l}$ to match the transpose of the feedback weights $\mathbf{B}_{l}$ roughly during training. This makes the error teaching signal (5) similar to backpropagation (3), thus enabling the network to learn the task.
3 Random noise pretraining with feedback alignment
Algorithm 1 Random noise pretraining
1:procedure Random noise pretraining($network$ $f_{\theta}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{d}$)
2:     for each epoch do
3:         for each batch do
4:              $\mathbf{x}\sim\mathcal{N}(\mu,\sigma^{2}),\;\mathbf{y}\sim\mathcal{U}(0,d-1)$triangleright sampling random noise
5:              $\mathcal{L}(\theta)=\text{Loss}(f_{\theta}(\mathbf{x}),\mathbf{y})$triangleright forward pass, equation (1), (2)
6:              $\delta_{L}=\frac{\partial\mathcal{L}}{\partial\mathbf{o}_{L}}$triangleright compute error
7:              for layer $l$ = $L$-1 to 00 do
8:                  $\mathbf{W}_{l}=\mathbf{W}_{l}-\eta\,\boldsymbol{\delta}_{l+1}\,\mathbf{h}_{l}^{T}$triangleright update weights, equation (4)
9:                  $\boldsymbol{\delta}_{l}=(\mathbf{B}_{l}\,\boldsymbol{\delta}_{l+1})\odot\phi^{\prime}(\mathbf{o}_{l})$triangleright compute error, equation (5)
@@ 3 Random noise pretraining with feedback alignment
During the developmental stage, spontaneous neural activity in the brain plays a critical role in shaping and refining neural circuits. Initially wired immature neural circuits undergo modifications of their connections through the processes of regulated cell formation, apoptosis, and synapse refinement through spontaneous neural activity [4, 5, 6, 7, 8]. These pre-sensory activities and development processes are universal across sensory modalities, such as the visual [46, 47], auditory [48, 49], and sensorimotor systems [50, 51]. We focus here on a few characteristics of spontaneous neural activity in the brain. Spontaneous neural activity is not correlated to external stimuli but can refine and optimize neural circuits, before interaction with the external world can take place.
@@
Here, we propose a type of random training that is inspired by the spontaneous and prenatal neural activity in the brain to pretrain the neural network (Algorithm 1). In every iteration, we sampled random noise inputs $\mathbf{x}$ from a Gaussian distribution $\mathcal{N}(0,I)$ and random labels $\mathbf{y}$ from a discrete uniform distribution $\mathcal{U}(0,N_{\text{readout}}-1)$, without any correlation. The network $f_{\theta}$ was initialized with random weights and trained with the feedback alignment algorithm. In this study, we examined the effects of random noise training on the subsequent learning processes in model neural networks to understand the potential benefits of pretraining with random noise in biological brains and whether this strategy is applicable to machine learning algorithms.
@@ 4.1 Weight alignment to synaptic feedback during random noise training
To simulate a neural network initially wired by random weights and fixed random synaptic feedback, we adopted a network setting from the feedback alignment algorithm (Figure 1a) in which the weight transport problem can be avoided through the use of fixed random synaptic feedback. Thus, unlike backpropagation, this process is considered possible to exist in biological neural networks with local synaptic connections (Figure 1b). We used a two-layer feedforward neural network with ReLU nonlinearity for classification, $f_{\theta}:\mathbb{R}^{784}\rightarrow\mathbb{R}^{10}$ with 100 neurons in the hidden layer. By means of random noise training (Algorithm 1), we trained the neural network with random inputs sampled from a Gaussian distribution $\mathcal{N}(0,I)$, with labels also randomly sampled independently (Figure 1c).
@@
We observed that the training loss decreased noticeably during random training, even in the absence of meaningful data and even when $\mathbf{x}$ and $\mathbf{y}$ are randomly paired (Figure 1d). During the random training process, we focused on the alignment between the forward weights and the synaptic feedback. As described in the literature [32], the alignment of $\mathbf{W}_{l}$ and $\mathbf{B}_{l}$, i.e., similarity between $\delta_{\text{BP}}$ and $\delta_{\text{FA}}$, is crucial for calculating the error teaching signal precisely. To evaluate the alignment, we used cosine similarity, which is widely used for measuring the distance between two vectors.
@@
Definition. Given the forward weights $\mathbf{W}_{l}\in\mathbb{R}^{m\times n}$ and backward weights $\mathbf{B}_{l}\in\mathbb{R}^{n\times m}$, we measured alignment using cosine similarity. We define that $\mathbf{W}_{l}$ and $\mathbf{B}_{l}$ as aligned if the angle $\angle(\mathbf{W}_{l}^{T})_{i},(\mathbf{B}_{l})_{i}$ is significantly smaller than 90 degrees.
@@
Notably, we found that the weights of neurons are aligned to the corresponding synaptic feedback weights during the random training process (Figure 1e). We also observed that the angle between the forward weights and synaptic feedback of individual neurons in the hidden layer decreased asymptotically during random training. In a randomly initialized network, the alignment angle appeared to be close to 90°, demonstrating that the backward error signal is randomly distributed (Figure 1e, left, alignment angle in an untrained network vs. 90°, $n=100$, one-sample t-test, NS, $P=0.492$). However, after random training, the alignment angle decreased significantly, implying that the backward teaching signal becomes valid to back-propagate errors (Figure 1e, right, alignment angle in an untrained network vs. a randomly trained network, $n=100$, two-sample t-test, ${}^{*}P<0.001$). We confirmed that this is not simply due to input bias under a particular condition but is reproduced robustly with various input conditions (Figure 1f). These results suggest that neural networks can pre-learn how to back-propagate errors through random noise training.